# Disentangling $\gamma-\beta$: the 4th-order velocity moments based on spherical Jeans analysis

### v1

@@ Abstract
Distinguishing a core and a cusp within dark matter halos is complexified by the existence of mass-anisotropy degeneracy, where various combinations of velocity anisotropy ($\beta$) and inner density slope ($\gamma$) yield similar observational signatures. We construct a dynamical model that incorporates the 4th-order velocity moments to alleviate this challenge. The inclusion of the 4th-order velocity moments enables stars’ line-of-sight velocity distribution (LOSVD) to be flexible. This flexible LOSVD can cover from a thin-tailed to a heavy-tailed distribution that is inaccessible if only the 2nd-order moments are considered. We test our dynamical model using mock galaxies and find that a ratio of the global line-of-sight velocity dispersion of the mock galaxy to the velocity error measurement $\sigma_{\rm los,global}/\Delta v_{\rm los}\gtrsim 4$ is required to avoid obtaining systematically biased results. This bias arises from the strong dependency of the 4th-order moments on the LOSVD tails, and not even increasing the sample size to $10^{4}$ stars can mitigate its effect. In that velocity ratio, $\beta$ is recovered within $\sim 1\sigma$ even when the sample size is only 500 stars, regardless of the recovery of the other parameters. However, the estimation of $\gamma$ varies, depending on the degree to which LOSVDs deviate from Gaussianity. Because of the more significant change in its LOSVDs, a cored dark halo is more likely to be identified than a cusp.
@@ 1 Introduction
The $\Lambda$CDM theory has been facing some challenges, specifically in the subgalactic scales, with the core-cusp problem (Flores & Primack, 1994; Moore, 1994; Burkert, 1995) being one of the oldest (see Bullock & Boylan-Kolchin (2017) for a review). This problem signifies a discrepancy in the central density slope of dark matter halos as inferred from simulations involving only dark matter particles compared to those inferred from the analysis of observational data. Dark matter-only simulations predict the universality of the dark matter density profile that rises steeply in the inner part following $\rho(r)\propto r^{-\gamma}$, with $\gamma\approx 1.0$ regardless the initial condition, well known as the NFW profile (Navarro et al., 1996). This universal property is in tension with the results from kinematics analyses using observational data from various types of dwarf galaxies that do not converge and instead favor a diverse inner density slope where $0\lesssim\gamma\lesssim 1.5$ (e.g., Oh et al. (2015)).
@@
While the existence of the core profile remains a subject of debate and is potentially constrained by modeling limitations (de Blok, 2010), the focus of probing a dark matter density profile has been redirected from gas-rich dwarf spirals to the nearby region; satellites of the Milky Way. The Milky Way dwarf spheroidal (dSph) and ultra-faint dwarf (UFD) satellites are among the most dark matter-dominated as well as the oldest gravitational systems known at the moment (Simon, 2019; Battaglia & Nipoti, 2022). These types of galaxies are believed to be the building blocks of the MW-sized galaxies and retain the most pristine relics of ancient structures from the early universe. Their unbeatable low fraction of baryon content, even inside the half-light radius, makes them the most attractive sites to test dark matter theories because the baryonic effects are less significant compared to galaxies with a higher fraction of baryonic component. Their proximity also offers a key advantage that enables their member stars to be resolved, allowing for the line-of-sight velocity, projected position, and metallicity of the member stars to be obtained individually.
@@
Modeling the internal dynamics of dSphs and UFDs has been considered one of the primary pathways in addressing the core-cusp problem (see Battaglia et al. (2013) for a review). Despite the advantages mentioned earlier, distinguishing between a core and a cusp remains challenging. Obtained results are often unable to robustly discriminate the two profiles from each other based on the available data set (Breddels et al., 2013; Geringer-Sameth et al., 2015). One central problem is the mass-anisotropy degeneracy, i.e., effects due to only three phase spaces (the position on the sky plane and line-of-sight velocity) that are observable. Consequently, there exists a broad range of models with velocity anisotropies and density profiles consistent with observations. Improvements from various directions are made to break this degeneracy.
@@
Where available, the use of more than one chemodynamically distinct stellar population can partially alleviate the degeneracy (Walker & Peñarrubia, 2011; Amorisco & Evans, 2012; Kowalczyk & Łokas, 2022). Motivated by the elongated observed stellar distribution and non-sphericity of simulated dark halo, there are also attempts to explore axisymmetric models, investigating different kinematical signatures along the major and minor axes (Hayashi & Chiba, 2012, 2015; Hayashi et al., 2020). Other groups aim to mitigate this degeneracy by developing Schwarzschild orbital-based superposition methods, assuming either a spherically symmetric (Breddels et al., 2013; Kowalczyk et al., 2019) or axisymmetric (Jardel & Gebhardt, 2012; Jardel et al., 2013; Hagen et al., 2019) underlying gravitational potential. Another approach involves utilizing higher-order moments through the virial shape parameter (Richardson & Fairbairn, 2014; Genina et al., 2020) or Jeans modeling (Łokas, 2002, 2009; Richardson & Fairbairn, 2013).
@@
The limited exploration through the higher-order moments primarily stems from their heavy reliance on stars in the velocity tails and large sample size requirements (Merrifield & Kent, 1990). The analysis involving higher-order moments is further complicated by the substantial impact of interlopers, which can significantly alter the velocity distribution in the tails where the star count is low. Nevertheless, this situation is going to be improved in the near future, thanks to the Subaru Prime Focus Spectrograph project that will increase the kinematic sample in each of the targeted MW’s dwarf galaxies to approximately 5,000 stars with expected velocity error measurement of $\sim$2 km s-1 (Takada et al., 2014; Tamura et al., 2016; Hayashi et al., 2023a). Therefore, exploring the ability and behavior of the higher-order moments to mitigate mass-anisotropy degeneracy becomes a subject of investigation. In this work, we construct a dynamical model based on the spherically symmetric 2nd-order combined with 4th-order Jeans equations.
@@
This paper is organized as follows. We initiate by detailing the dynamical model, which relies on the spherically symmetric 2nd-order and 4th-order Jeans equations, in Section 2. Section 3 presents the outcomes derived from applying the model to mock data. In section 4, we discuss the dependency on the choice of prior, the recovery of $\gamma$, and the recovery of $\beta$. We summarize our main conclusions in section 5.
@@ 2 Models
This section consists of an explanation of the dark matter and star distribution we adopt (Section 2.1), the 2nd-order (Section 2.2) and 4th-order (Section 2.3) Jeans modeling that we apply, the fitting procedure to marginalize parameters (Section 2.4), and details of the mock data (Section 2.5).
@@ 2.1 Dark matter density and light profiles
In Equation (1), $\rho_{0}$ is the dark matter scale density, $a_{\rm dm}$ is the dark matter scale length, $\gamma$ is the inner density slope, $\beta_{\rm trans}$ is the transition sharpness between the inner and the outer slope density profile. A smaller value of $\beta_{\rm trans}$ means a smoother, more gradual change from the inner slope to the outer slope. The last parameter, $\alpha$, is the outer slope density profile that dominates the slope of the density profile in the galactocentric distance larger than $a_{\rm dm}$. This model provides a wide range of density profiles, including the cored profile ($\gamma\approx 0$), the cuspy profile ($\gamma\approx 1$), and any possible profiles between the two.
@@ 2.2 Spherically symmetric 2nd-order Jeans equations
Assuming that a gravitational system is static and spherical, the motion of its stars moving under the influence of underlying gravitational potential $\Phi(r)$ can be explained by the distribution function $f(\textbf{r},\textbf{v})$. However, the distribution function is inaccessible in the real application. Jeans equations link the inaccessible distribution function to the gravitational potential via velocity moments. Defining velocity moments with spherical coordinates ($r$, $\theta$, $\phi$) as
@@ 2.3 Spherically symmetric 4th-order Jeans equations
In the unbinned 2nd-order only Jeans modeling, one usually links the theoretical l.o.s. velocity dispersion $\sigma_{\rm los}$ to the LOSVD by assuming that its distribution is Gaussian. Together with the velocity error measurement $\Delta v_{\rm los}$, which has a Gaussian distribution as well, these two quantities characterize the LOSVD for each member star at position $R$. However, taking into account the 4th-order moments, we need to provide more freedom to the distribution function to change its shape. For this purpose, we adopt two kernels (uniform kernel and Laplacian kernel, see Appendix for details) explained in Sanders & Evans (2020) to link the theoretical l.o.s. velocity dispersion and kurtosis to the LOSVD. Figure 1 shows how the value of kurtosis changes the shape of a distribution in this model. Another notable advantage of adopting these kernels is that the velocity error measurements for each star can be convolved into the distribution. Therefore, stars with precise measurements do not lose their strong signature of kinematical information because of averaging the error or their l.o.s. velocity is unnecessary.
Figure 1: The changing shape of LOSVDs for different choices of $\kappa_{\rm los}$.
@@ 2.4 Fitting Procedure
In order to understand to what extent this model can recover the free parameters, we explore their most likely value by estimating the probability of each observed star via the calculated LOSVD, which is characterized by theoretical l.o.s. velocity dispersion and kurtosis. This work avoids a binned analysis because the results suffer from a strong dependency on the choice of binning and, hence, are prone to bias. The binned analysis is also avoided because extracting the observed $\kappa_{\rm los}$ by binning the sample can introduce a potentially large error of $\gtrsim 20\%$ unless each bin contains $\gtrsim 150$ stars (Merrifield & Kent, 1990).
We assume that there is no net streaming motion, therefore the LOSVD is symmetrically centered at the systemic velocity of the stellar system $v_{\rm sys}$. For $N$ total number of sample stars in a stellar system, $v_{{\rm los,}i}\pm\delta_{i}$ denoting the observed l.o.s. velocity and its uncertainty for the $i$-th star, the likelihood function is given by
$L=\prod_{i=1}^{N}\frac{1}{(\delta_{i}^{2}+\sigma_{{\rm los,}i}^{2})^{1/2}}f_{s}(w_{i}),$ (12)
@@
$w_{i}^{2}=\frac{(v_{{\rm los,}i}-v_{\rm sys})^{2}}{\delta^{2}_{i}+\sigma_{{\rm los,}i}^{2}},$ (13)
and the choice of $f_{s}(w)$ is determined by the value of $\kappa_{\rm los}$. We use the Laplacian kernel if $\kappa_{\rm los}>3$, the uniform kernel if $\kappa_{\rm los}<3$, and Gaussian kernel if $\kappa_{\rm los}=3$. These choices of kernel allow the shape of LOSVDs to be flexible, covering more thin-tailed to heavy-tailed distributions than a Gaussian, with the limitation being the pure uniform ($\kappa_{\rm los}=1.8$) and pure Laplacian ($\kappa_{\rm los}=6.0$) distribution at each extreme case. The theoretical velocity dispersion for the $i$-th star $\sigma_{{\rm los,}i}$ is calculated through Equation 7 at the position $R_{i}$. A flat or log-flat prior is applied to marginalize the free parameters in these given ranges:
@@
Even though it is unphysical to have a system with $\gamma<0$, which means a ’hole’ in the center, we do not adopt a more restrictive prior range on $\gamma$ that excludes the negative region. The detail concerning this selection is explained in Section 4.2.
Finally, we apply the Markov Chain Monte Carlo in the frame of the Metropolis-Hastings algorithm (Metropolis et al., 1953; Hastings, 1970) to obtain the posterior distribution of the free parameters. To minimize the effects of initial conditions, we perform a burn-in process on the obtained samples.
@@ 2.5 Mock data
To test our constructed dynamical modeling, we generate mock data samples using StarSampler [^1] https://github.com/maoshenl/StarSampler (Liu, 2019). For a dark matter potential, we set two different density profiles: one is an NFW profile $(\alpha,\beta_{\mathrm{trans}},\gamma)=(3,1,1)$, and another is a cored NFW profile $(\alpha,\beta_{\mathrm{trans}},\gamma)=(3,1,0)$. The central density $\rho_{0}$ is adopted as $0.064\times 10^{9}$ [$M_{\odot}$ pc-3] for each case of DM potential.
For a light profile, we can model the stellar density profile as a Hernquist/Zhao profile (Equation 1). In this paper, we generate mock stars following a Plummer profile in a specific case of the Hernquist/Zhao profile. To this end, we set the parameters as $(\alpha_{\ast},\beta_{\mathrm{trans},\ast},\gamma_{\ast})=(5,2,0)$, where $\alpha_{\ast},\beta_{\mathrm{trans},\ast},\gamma_{\ast}$ are the outer slope, transition sharpness, and inner slope of stellar density profile, respectively. Since the stars are embedded centrally in their respective DM halos, the parameter $b=a_{\mathrm{dm}}/a_{\ast}$ is adopted as 10. The details of the mock data are summarized in Table 1.
Table 1: Input parameters of the mock data
Parameter Input value in Mock Galaxy 01 Input value in Mock Galaxy 02
$\beta$ 0.0 0.0
$\alpha$ 3.0 3.0
$\beta_{\rm trans}$ 1.0 1.0
$\gamma$ 1.0 0.0
$\rho_{0}$/[M/pc3] $0.064\times 10^{9}$ $0.064\times 10^{9}$
$b$ 10 10
@@
Figure 2 shows the intrinsic l.o.s. velocity dispersion (left panel) and kurtosis (right panel) profile of the mock data set we generate. Turquoise lines denote the profiles for the cuspy Mock Galaxy 01, while gold lines denote the cored Mock Galaxy 02. Note from the point of view of the global stellar l.o.s. velocity dispersion,
$\sigma_{\rm los,global}=\sqrt{\sum_{i=1}^{N}\frac{(v_{{\rm los},i}-v_{\rm sys})^{2}}{N-1}},$ (14)
where $N$ is the number of sample stars. Mock Galaxy 01 global l.o.s. velocity dispersion of $\sigma_{\rm los,global}=8.6$ km s-1 lies in the range that is similar to the ”classical” MW dSphs, e.g., Draco ($\sigma_{\rm los,global}=9.0^{+0.3}_{-0.3}$ km s-1), Fornax ($\sigma_{\rm los,global}=11.4^{+0.4}_{-0.4}$ km s-1), Sextans ($\sigma_{\rm los,global}=8.4^{+0.4}_{-0.4}$ km s-1), Ursa Minor ($\sigma_{\rm los,global}=8.0^{+0.3}_{-0.3}$ km s-1). Mock Galaxy 02 with $\sigma_{\rm los,global}=3.3$ km s-1, on the other hand, resembles the less massive category of an ultra-faint dwarf galaxy, e.g., BootesI ($\sigma_{\rm los,global}=2.4^{+0.9}_{-0.5}$ km s-1), Coma Berenices ($\sigma_{\rm los,global}=4.6^{+0.8}_{-0.8}$ km s-1), CraterII ($\sigma_{\rm los,global}=2.7^{+0.3}_{-0.3}$ km s-1), SegueI ($\sigma_{\rm los,global}=3.7^{+1.4}_{-1.1}$ km s-1) (Battaglia et al., 2022). Because of its divergent density profile in the center, $\sigma_{\rm los}(R)$ of Mock Galaxy 01 experiences a deflection increase at some point toward the center. This is not observed in the cored Mock Galaxy 02 due to its constant and fairly lower density toward the center.
Figure 2: The left panel shows the intrinsic l.o.s. velocity dispersion as a function of the projected radius. The right panel shows the intrinsic l.o.s. kurtosis as a function of the projected radius. Each turquoise (gold) line represents the profiles of Mock Galaxy 01 (02).
The difference of $\gamma$ between two mock galaxies gives the difference in the l.o.s. kurtosis profile. Despite the outer regions being relatively similar due to identical $\alpha$, the inner regions exhibit a continuous divergence in $\kappa_{\rm los}(R)$ towards the center. While $\kappa_{\rm los}(R)$ for Mock Galaxy 01 merely departs from Gaussianity $\kappa_{\rm los}=3$ throughout its extent, $\kappa_{\rm los}(R)$ for Mock Galaxy 02 significantly departs from Gaussianity toward the center, reaching $\kappa_{\rm los}\approx 6$ in the centermost part. The large $\kappa_{\rm los}$ indicates strong tails in high $v_{\rm los}$ regimes and a spiky peak around $v_{\rm los}=v_{\rm sys}$ of the LOSVD.

### v2

@@ Abstract
Distinguishing a core and a cusp within dark matter halos is complexified by the existence of mass-anisotropy degeneracy, where various combinations of velocity anisotropy ($\beta$) and inner density slope ($\gamma$) yield similar observational signatures. We construct a dynamical model that incorporates the 4th-order velocity moments to alleviate this challenge. The inclusion of the 4th-order velocity moments enables stars’ line-of-sight velocity distribution (LOSVD) to be flexible. This flexible LOSVD can cover from a thin-tailed to a heavy-tailed distribution that is inaccessible if only the 2nd-order moments are considered. We test the model on four mock galaxies having isotropic orbits, $\beta=0$: two resembling dwarf spheroidal galaxies (dSphs) and two resembling ultra-faint dwarfs (UFDs) in terms of velocity dispersion. Each category includes one galaxy with a cuspy NFW profile and one with a cored density profile. Results show that a ratio of the global velocity dispersion to velocity error, $\sigma_{\rm los,global}/\delta v_{\rm los}\gtrsim 4$, is crucial to avoid systematic biases arising from the strong sensitivity of 4th-order moments to the LOSVD tails. In cases where this velocity ratio condition is met, our model reliably recovers $\gamma$ in dSph mock galaxies, with the true value recovered within $\sim 1\sigma$, and strongly excludes a cuspy NFW profile for the cored dSph mock galaxy. However, recovering the density profiles of UFDs remains challenging due to their intrinsically low velocity dispersions.
@@ 1 Introduction
The $\Lambda$CDM theory has been facing some challenges, specifically in the subgalactic scales, with the core-cusp problem (Flores & Primack, 1994; Moore, 1994; Burkert, 1995) being one of the oldest (see Bullock & Boylan-Kolchin (2017) and Salucci (2019) for reviews). This problem signifies a discrepancy in the central density slope of dark matter halos as inferred from simulations involving only dark matter particles compared to those inferred from the analysis of observational data. Dark matter-only simulations predict the universality of the dark matter density profile that rises steeply in the inner part following $\rho(r)\propto r^{-\gamma}$, with $\gamma\approx 1.0$ regardless the initial condition, well known as the NFW profile (Navarro et al., 1996). This universal property is in tension with the results from kinematics analyses using observational data from various types of dwarf galaxies that do not converge and, instead, favor a diverse inner density slope where $0\lesssim\gamma\lesssim 1.5$ (e.g., Oh et al. (2015)).
@@
While the existence of the core profile remains a subject of debate and is potentially constrained by modeling limitations (de Blok, 2010), the focus of probing a dark matter density profile has been redirected from gas-rich dwarf spirals (Burkert, 1995; Salucci & Burkert, 2000) to the nearby region; satellites of the Milky Way. The Milky Way dwarf spheroidal (dSph) and ultra-faint dwarf (UFD) satellites are among the most dark matter-dominated galaxies known at the moment (Simon, 2019; Battaglia & Nipoti, 2022). These types of galaxies are believed to be the building blocks of the MW-sized galaxies and retain the most pristine relics of ancient structures from the early universe. Their unbeatable low fraction of baryon content, even inside the half-light radius, makes them the most attractive sites to test dark matter theories because the baryonic effects are less significant compared to galaxies with a higher fraction of baryonic component. Their proximity also offers a key advantage that enables their member stars to be resolved, allowing for the line-of-sight velocity, projected position, and metallicity of the member stars to be obtained individually.
@@
Modeling the internal dynamics of dSphs and UFDs has been considered one of the primary pathways in addressing the core-cusp problem (see Battaglia et al. (2013) for a review). Despite the advantages mentioned earlier, distinguishing between a core and a cusp remains challenging. Obtained results are often unable to robustly discriminate the two profiles from each other based on the available data set (Breddels et al., 2013; Geringer-Sameth et al., 2015). One central problem is the mass-anisotropy degeneracy, where a wide range of models with varying velocity anisotropies and density profiles can reproduce the observed kinematics. Improvements from various directions are made to break this degeneracy.
@@
Where available, the use of more than one chemodynamically distinct stellar population can partially alleviate the degeneracy (Walker & Peñarrubia, 2011; Amorisco & Evans, 2012a; Kowalczyk & Łokas, 2022). Motivated by the elongated observed stellar distribution and non-sphericity of simulated dark halo, there are also attempts to explore axisymmetric models, investigating different kinematical signatures along the major and minor axes (Hayashi & Chiba, 2012, 2015; Hayashi et al., 2020). Other groups aim to mitigate this degeneracy by developing Schwarzschild orbital-based superposition methods, assuming either a spherically symmetric (Breddels et al., 2013; Kowalczyk et al., 2019) or axisymmetric (Jardel & Gebhardt, 2012; Jardel et al., 2013; Hagen et al., 2019) underlying gravitational potential. In addition, methods based on distribution functions (DFs) offer a robust framework for dynamical modeling. These DF-based methods, as developed by Amorisco & Evans (2011), Pascale et al. (2018), and Strigari et al. (2017), ensure self-consistency by directly modeling the phase-space distribution of stars. By satisfying the collisionless Boltzmann equation, these methods naturally reproduce the observed line-of-sight velocity distributions (LOSVDs). Another approach involves utilizing higher-order moments through the virial shape parameter (Richardson & Fairbairn, 2014; Genina et al., 2020) or Jeans modeling (Łokas, 2002, 2009; Richardson & Fairbairn, 2013).
@@
The limited exploration through the higher-order moments primarily stems from their heavy reliance on stars in the velocity tails and large sample size requirements (Merrifield & Kent, 1990). The analysis involving higher-order moments is further complicated by the substantial impact of interlopers, which can significantly alter the velocity distribution in the tails where the star count is low. Nevertheless, this situation is going to be improved in the near future, thanks to the Subaru Prime Focus Spectrograph project that will increase the kinematic sample in each of the targeted MW’s dwarf galaxies to approximately 5,000 stars with expected velocity error measurement of $\sim$2 km s-1 (Takada et al., 2014; Tamura et al., 2016; Hayashi et al., 2023). Therefore, exploring the ability and behavior of the higher-order moments to mitigate mass-anisotropy degeneracy becomes a subject of investigation. In this work, we construct a dynamical model based on the spherically symmetric 2nd-order combined with 4th-order Jeans equations.
@@
This paper is organized as follows. We initiate by detailing the dynamical model, which relies on the spherically symmetric 2nd-order and 4th-order Jeans equations, in Section 2. Section 3 presents the outcomes derived from applying the model to mock data. In section 4, we discuss the recovery of $\gamma$ and the recovery of $\beta$. We summarize our main conclusions in section 5.
@@ 2 Models
In this section, we provide a detailed overview of the components and methods used in our analysis. First, we describe the adopted distributions for dark matter and stars (Section 2.1). Next, we outline the 2nd-order (Section 2.2) and 4th-order (Section 2.3) Jeans modeling frameworks employed in this study. We then explain the parameter marginalization process in the fitting procedure (Section 2.4), followed by a description of the mock data used for testing and validation (Section 2.5).
@@ 2.1 Dark matter density and light profiles
In Equation (1), $\rho_{0}$ is the dark matter scale density, $a_{\rm dm}$ is the dark matter scale length, $\gamma$ is the inner density slope, and $\beta_{\rm trans}$ is the transition sharpness between the inner and the outer slope density profile. A smaller value of $\beta_{\rm trans}$ means a smoother, more gradual change from the inner slope to the outer slope. The last parameter, $\alpha$, is the outer slope density profile that dominates the slope of the density profile in the galactocentric distance larger than $a_{\rm dm}$. This model provides a wide range of density profiles, including the cored profile ($\gamma\approx 0$), the cuspy profile ($\gamma\approx 1$), and any possible profiles between the two.
@@ 2.2 Spherically symmetric 2nd-order Jeans equations
Assuming that a gravitational system is static and spherical, the motion of its stars moving under the influence of underlying gravitational potential $\Phi(r)$ can be explained by the distribution function $f(\textbf{r},\textbf{v})$. However, the distribution function is inaccessible in the real application. Jeans equations link the inaccessible distribution function to the gravitational potential via velocity moments. We define the velocity moments in spherical coordinates ($r,\theta,\phi$) as:
@@ 2.3 Spherically symmetric 4th-order Jeans equations
In the unbinned 2nd-order only Jeans modeling, one usually links the theoretical l.o.s. velocity dispersion $\sigma_{\rm los}$ to the LOSVD by assuming that its distribution is Gaussian. Together with the velocity error measurement $\delta v_{\rm los}$, which has a Gaussian distribution as well, these two quantities characterize the LOSVD for each member star at position $R$. However, taking into account the 4th-order moments, we need to provide more freedom to the proposed LOSVD to change its shape. For this purpose, we adopt two kernels (uniform and Laplacian kernels, see Appendix A for details) explained in Sanders & Evans (2020) to link the theoretical l.o.s. velocity dispersion and kurtosis to the LOSVD. Figure 1 shows how the value of kurtosis changes the shape of a distribution in this model. Another notable advantage of adopting these kernels is that the velocity error measurements for each star can be convolved analytically into the distribution. Therefore, stars with precise measurements do not lose their strong signature of kinematical information because of averaging the error or their l.o.s. velocity is unnecessary.
Figure 1: Probability density functions based on uniform and Laplacian kernels for scaled line-of-sight velocities across various choices of $\kappa_{\rm los}$.
@@ 2.4 Fitting Procedure
This work avoids a binned analysis because the results suffer from a strong dependency on the choice of binning and, hence, are prone to bias. The binned analysis is also avoided because extracting the observed $\kappa_{\rm los}$ by binning the sample can introduce a potentially large error of $\gtrsim 20\%$ unless each bin contains $\gtrsim 150$ stars (Merrifield & Kent, 1990). Instead of employing binning to estimate the most likely values of the free parameters, we adopt a different approach. Specifically, we calculate the probability of each star having the l.o.s. velocity $v_{{\rm los,}i}$, given a set of parameters $\theta(\beta,\alpha,\beta_{\rm trans},\gamma,\rho_{0},b)$ that describe the LOSVD at the star’s projected position $R_{i}$. The LOSVD’s width is characterized by the l.o.s velocity dispersion $\sigma_{{\rm los,}i}(R_{i}|\theta)$ calculated through Equation (7) and convolved with the observational velocity error $\delta v_{\mathrm{los,}i}$. Meanwhile, the l.o.s kurtosis $\kappa_{\rm los}(R_{i}|\theta)$ that is calculated through Equation (11) controls the LOSVD’s shape.
We assume that there is no net streaming motion. Therefore, LOSVDs are symmetrically centered at the systemic velocity of the stellar system $v_{\rm sys}$. For $N$ total number of sample stars in a stellar system, the likelihood function is given by
$L=\prod_{i=1}^{N}\frac{1}{(\delta v_{\mathrm{los,}i}^{2}+\sigma_{{\rm los,}i}^{2})^{1/2}}f_{s}(w_{i}),$ (12)
@@
$w_{i}^{2}=\frac{(v_{{\rm los,}i}-v_{\rm sys})^{2}}{\delta v_{\mathrm{los,}i}^{2}+\sigma_{{\rm los,}i}^{2}},$ (13)
and the choice of $f_{s}(w)$ is determined by the value of $\kappa_{\rm los}$. If a combination of parameters results in $\kappa_{\rm los}<3$, the uniform kernel is used, and if $\kappa_{\rm los}>3$, the Laplacian kernel is applied. Additionally, $\kappa_{\rm los}$ is linked to the uniform kernel via Equations (A2) and (A6), and to the Laplacian kernel via Equations (A9) and (A10). These equations ensure that for any given $\kappa_{\rm los}$, there exists a unique solution for the distribution $f_{s}(w)$. These choices of kernels allow the shape of LOSVDs to be flexible, covering more thin-tailed to heavy-tailed distributions than a Gaussian, with the limitation being the pure uniform ($\kappa_{\rm los}=1.8$) and pure Laplacian ($\kappa_{\rm los}=6.0$) distribution at each extreme case. Detailed properties of these kernels are given in Appendix A. A flat or log-flat prior is applied to marginalize the free parameters in these given ranges:
@@
Even though it is unphysical to have a system with $\gamma<0$, which means a ’hole’ in the center, we do not adopt a more restrictive prior range on $\gamma$ that excludes the negative region. The detail concerning this selection is explained in Section B. For simplicity, we assume that the projected stellar distribution is well-estimated from photometric measurements. Therefore, no parameter describing the stellar distribution is treated as a free parameter in our analysis. An alternative approach that might be more accommodating for observational uncertainties in the stellar distribution is demonstrated by Chang & Necib (2021), where a two-layered fitting procedure is employed: the first step involves estimating the stellar distribution, followed by a second step to constrain the gravitational potential.
Finally, we use the Markov Chain Monte Carlo (MCMC) method within the Metropolis-Hastings framework (Metropolis et al., 1953; Hastings, 1970) to derive the posterior distribution of the free parameters. For this analysis, we employ a custom-made code specifically developed for this work. Convergence of the MCMC chains is assessed through visual inspection. Our MCMC setup involves 20 walkers, each generating 25,000 samples for each realization, with the first 5,000 samples discarded in the burn-in process to eliminate the initial point effects. After post-processing, this results in a total of 400,000 samples for each realization. Lastly, posterior distributions are generated by averaging across all realizations within the same scenario.
@@ 2.5 Mock data
We generate mock galaxies using StarSampler [^1] https://github.com/maoshenl/StarSampler (Liu, 2019) to apply and test our constructed dynamical model. StarSampler is a Python-based tool developed to create random samples from any stationary distribution function (DF) specified by the user. This DF represents the probability density of stellar positions and velocities within six-dimensional phase space. The module is versatile, enabling the sampling of DFs that describe stellar systems, whether they are self-gravitating or situated within an external potential created by other sources, which is a dark matter halo in this case. Two sampling techniques are available, allowing users to select based on their desired level of efficiency and accuracy: the rejection sampling method, which delivers exact results but may be computationally inefficient, and the importance sampling method, which is faster but provides an approximate solution.
In this study, we set all mock galaxies to be isotropic, $\beta=0$, for the sake of simplicity. The light profile of the galaxies is modeled using a Plummer profile, which is a specific case of the Generalized Hernquist Model characterized by $(\alpha_{\ast},\beta_{\mathrm{trans},\ast},\gamma_{\ast})=(5,2,0)$. Here, $\alpha_{\ast}$, $\beta_{\mathrm{trans},\ast}$, and $\gamma_{\ast}$ represent the outer slope, transition sharpness, and inner slope of the stellar density profile, respectively. For the dark matter potential, we adopt two distinct density profiles: the first is an NFW profile defined by $(\alpha,\beta_{\mathrm{trans}},\gamma)=(3,1,1)$, and the second is a cored NFW profile with $(\alpha,\beta_{\mathrm{trans}},\gamma)=(3,1,0)$. There are two galaxies generated for each density profile to cover two different regimes: dSphs and the less bright, and less massive UFDs (Simon, 2019; Battaglia et al., 2022). This is done by differing the input value of dark matter scale density $\rho_{0}$ and dark matter scale length $a_{\rm dm}$ while keeping the intrinsic velocity dispersion profiles to be similar. All details of the mock galaxies are summarized in Table 1.
Table 1: Input parameters of the mock data.
Parameter Cored dSph Cuspy dSph Cored UFD Cuspy UFD
$\beta$ 0.0 0.0 0.0 0.0
$\alpha$ 3.0 3.0 3.0 3.0
$\beta_{\rm trans}$ 1.0 1.0 1.0 1.0
$\gamma$ 0.0 1.0 0.0 1.0
$\rho_{0}$/[M pc-3] $0.6$ $0.064$ $0.064$ $0.002$
$b$ 4.5 10 10 50
@@
Figure 2 shows the intrinsic density profile (left panel), l.o.s. velocity dispersion (middle panel), and l.o.s. kurtosis (right panel) profile of the mock galaxies. Solid lines denote cuspy mock galaxy profiles, while dash-dotted lines denote the cored one. Thick lines mark mock galaxies resembling the classical MW’s satellites, and the thinner lines are for mock galaxies similar to UFD’s.
Figure 2: Density, l.o.s. velocity dispersion, and l.o.s. kurtosis profiles for the mock galaxies are shown in the left, middle, and right panels, respectively. The purple dash-dotted lines represent galaxies with a cored density profile, while the orange solid lines indicate galaxies with a cuspy density profile. Mock galaxies resembling dSphs are shown with thicker lines, and those resembling UFDs are shown with thinner lines.
To ensure robust statistical analysis, we generate 10 realizations for each scenario set. These realizations are created for data sets containing 500 and 5,000 stars, respectively [^2] The velocity dispersion calculated using 5,000 stars indicates that the dSph mock galaxies resemble Sextans and Leo II, while the UFD mock galaxies are comparable to Segue I and Reticulum II (Battaglia et al., 2022). The choices of sample sizes are specifically designed to allow a comparison between the currently available data in dSphs and UFDs, and the expected data quality after the Subaru Prime Focus Spectrograph (PFS) survey (Takada et al., 2014). The independent results from each realization within the same scenario are then averaged to ensure consistency and mitigate the effects of statistical fluctuations. We assume two levels of measurement error in the line-of-sight velocity. Specifically, we set $\delta v_{\rm los}=2$ km s-1 to represent the typical measurement error in current datasets and $\delta v_{\rm los}=0.01$ km s-1 to suppress any biases associated with velocity uncertainties and identify any change present when we assign a larger one. For clarity and consistency, throughout this paper, we adopt the notation $R$ to represent the projected radius and $r$ to denote the three-dimensional galactocentric radius. This distinction ensures an unambiguous interpretation of results and parameters in the context of spherical symmetry.